OpenStudy (anonymous):

A triangular prism has a volume of 5 m3. If each dimension of the prism is tripled, what is the new volume of the prism? Answer 135 m3 45 m3 15 m3 405 m3 can i really just multiply the original volume by 3?

hero (hero):

Says "if each dimension is tripled" Key word: Dimension

OpenStudy (anonymous):

oh ok well then i may need some help on this...

OpenStudy (anonymous):

hmm maybe multiply by 3^3?

OpenStudy (anonymous):

u there?

OpenStudy (anonymous):

yes im here =) and

OpenStudy (anonymous):

how do i find the demensions thas what im gonna have a hard time finding

hero (hero):

You have to use the formula for volume of triangular prism

OpenStudy (anonymous):

mmk

OpenStudy (anonymous):

if not mistaken volume for triangular prism should be V=1/2 * bhl=5. b= base length, h= height, l= length between 2 triangles. for the new volume, Lets say A. A=1/2 * (3b)(3h)(3l)=(1/2)bhl(3^3)=5*27=135. so i think the answer would be 135.